OpenStudy (anonymous):

Solve 2x – 5 = 3(2x + 5). You must show all work

OpenStudy (anonymous):

im not really good at this but ill try

OpenStudy (anonymous):

2x-5=3(2x+5) -5 -5 0 -5 = -2 -.- i have no idea what im doing

OpenStudy (johnweldon1993):

Not quite @kittywhite here let me take a crack at it...

OpenStudy (johnweldon1993):

So you start off with 2x - 5 = 3(2x + 5) You first want to distribute that 3 into the parenthesis....distribute means multiply into each term...so what is 3 times 2x? what is 3 times 5? This will leave us with 2x - 5 = 6x + 15 what would be do next...?

OpenStudy (anonymous):

2x – 5 = 3(2x + 5) 2x - 5 = 6x + 15 2x -5 + 5 = 6x + 15 + 5 2x = 6x + 20 2x - 6x = 6x-6x + 20 -4x = 20 -4x/-4 = 20/-4 x = -5 I think..

OpenStudy (anonymous):

oh yeah i learned this in 6th grade i must have forgot 3x2x=6x? and 3x5=15

OpenStudy (johnweldon1993):

Very good @LoveYou*69 :)

OpenStudy (anonymous):

well im not that smart im only in 7th grade o.o well going to 7th

OpenStudy (anonymous):

I'm going into 10th, 9th right now, don't worry about it(: We all have those times.

OpenStudy (johnweldon1993):

I'm going into...umm *calculating* lol...14th XD

OpenStudy (anonymous):

So college :P?

OpenStudy (anonymous):

thank you @LoveYou*69

OpenStudy (johnweldon1993):

Haha yeah sophomore year :)

OpenStudy (johnweldon1993):

or no...2nd year lol

OpenStudy (anonymous):

Haha ooh, nice(:

OpenStudy (anonymous):

i feel like i have long term memory lost o.o i forgot how to use a x-axis y-axis graph with the negatives and positives

OpenStudy (anonymous):